# Tackling the Inherent Difficulty of Noise Filtering in RAG

### v1

@@ 6.2 Implementation Details
When calculating $\Delta{\bm{W}}$, we use Low Rank Adaption, which means we actually calculates $\Delta{\bm{W}}={\bm{A}}\cdot{\bm{B}}$, ${\bm{A}}\in\mathbb{R}^{h\times r}$, where $h$ is the hidden dimension and $r$ is the rank, in our experiments, we set $r=64$.
@@
We compare our method with LoRA mainly because our method focus on how to adjust the attention schema to make the model more robust to noise. But current research on robust RAG primarily focuses on either how to structure training data (RAG_robust; RAG_robust_finetune_25) or how to train models to better handle different types of noise. In contrast, our work addresses a different challenge: the standard self-attention mechanism inherently struggles to filter out noisy information. To mitigate this, we propose modifying the attention computation by incorporating a non-linear activation function to enhance robustness. Differential Transformer (differentialtransformer) also tries to adjust attention, but requires to train the model from scratch instead of fine-tuning based on existing model, and CrAM does not involve fine-tuning, so the comparison is unfair.
@@ 6.3 Main Results
Figure 3: The difference of attention score on answer tokens (mean(attn(answer))-mean(attn(other)).
Table 1 shows the performance when faced with explicit distracting documents, we evaluate two different setting where the query is placed behind the documents or ahead the documents. The result shows that with $g(x)$, the model can better distinguish between relevant and distracting documents, showing better performance. Also, we can observe that after fine-tuning, placing the query ahead of documents still helps, which further validate our theory in Section 5.2. Also, we can find that when setting $\xi=0$ performs similarly with LoRA fine-tuning, this is not surprising because they are technically equal, and they use the same number of parameters. The result shown in Table 1 might be high especially for 2wiki, this is mainly because we add gold documents to the context to show how can the model grab useful information from context. We also show the performance when all documents are retrieved in Table 4, it also shows that our method helps. We also conduct experiments when all parameters can be tuned and the result is shown in Appendix A
@@
We also conduct experiments on Qwen/Qwen2.5-7B-Instruct and mistralai/Mistral-7B-Instruct-v0.2, we show the performance in Table 2, which shows that our method also effectively helps the performance on Qwen and Mistral.
@@ 6.4 The Difference of Attention Score
By adding the activation function, our method can effectively distinguish between useful information and noise, to show this, we calculate the attention score gap between the tokens containing answer and other tokens (after softmax, mean(attn(answer))-mean(attn(other)), for clarity, we scale this gap by a factor of 1,000, the result is show in Figure 3. We do not show the result of ASQA because it is a long form QA, the answer is not directly indicated in the documents. The result shows that the rectification helps to recognize the answer and filter out noisy information.
6.5 Ablation Study
Figure 4: The performance when we set different activation function, $sum(\xi\cdot tanh(x),x)$ (sum), $\xi\cdot tanh(x)$ (tanh). We also show the performance when we use $g(x)=max(x,2\xi\cdot(sigmoid(x)-0.5))$ (sigmoid). $\xi$ is set to 3, and max means our method.
We also conduct experiments when $g(x)=\xi\cdot tanh(x)$ and $g(x)=\xi\cdot tanh(x)+x$. $g(x)=tanh(x)$ stands for the situation that $g(x)$ only focus on enlarge the margin between relevant and irrelevant tokens without further linear growth. And $g(x)=tanh(x)+x$ stands for the situation where the the steady growth process is missing, so the soft clamp will not work. We also try using $g(x)=max(x,\xi\cdot(sigmoid(x)-0.5))$ to show the performance when using sigmoid instead of tanh. As shown in Figure 4, simply use $tanh(x)$ or $tanh(x)+x$ has suboptimal performance and replacing tanhtanh with sigmoid(x)0.5sigmoid(x)-0.5 also helps the performance. We also discuss the efficiency of our method in Appendix A
@@ 6.5 Ablation Study
6.6 Performance with All Retrieved Documents
In previous experiments, we directly add the gold documents to the input to show that our method can effectively grab useful information, and with gold documents, we can easily locate the answer token for better analysis, this is also why the model performs good. Here we conduct experiments when all documents are retrieved by DPR, specifically, we use DPR to retrieve top 5 documents and use those retrieved documents as context. Then, we can observe from Table 4 that our method performs better for both 3 models and 5 datasets. We conduct experiments with the query placed ahead of documents.
@@ 7 Conclusion
In this paper, we highlight that noise filtering in RAG is inherently a difficult problem, judging is a token relevant may involve three or even more other tokens, and limited number of transformer layers can not effectively solve it. Then we show that simply fine-tuning the LLM may not be optimal as it will disturb the attention pattern, which could lead to worse performance. Then we propose a new fine-tuning method which can help to better filter out noise and take advantage of the relevant information, extensive experiments show that our method works well.
@@
The paper discusses the limitation of LLM when dealing with noisy information, showing that current LLMs can not effectively process noisy information. However, although a new fine-tuning method is proposed, it can not fully address the problem as it is a fine-tuning based on the trained model. It might help more if we train a model from scratch, but due to limited computational resource, this can only leave for furture work.
@@ A.2 Placing the Query Ahead Helps
Table 4: The performance when all documents are retrieved. Our method also shows better performance
@@ Appendix B Proofs
B.1 Proof of Theorem 4.1
@@
Theorem B.1.
if there exists $attn^{\prime}(x_{i},x_{j})={\bm{x}}_{i}({\bm{W}}+\Delta{\bm{W}}){\bm{x}}_{j}$ satisfying $attn^{\prime}(x_{i},x_{j})-\widehat{attn}(x_{i},x_{j})\leq\epsilon$, then we need
$\begin{split}\xi_{r}&\lesssim\ln\frac{1}{1-\epsilon}\\ c_{l}-c_{h}&\leq\xi_{r}+\ln\epsilon n\end{split}$
@@
B.2 MLP also fails to filter out noise

### v2

@@ 6.2 Implementation Details
When calculating $\Delta{\bm{W}}$, we use Low Rank Adaption, which means we actually calculates $\Delta{\bm{W}}={\bm{A}}\cdot{\bm{B}}$, ${\bm{A}}\in\mathbb{R}^{h\times r}$, where $h$ is the hidden dimension and $r$ is the rank, in our experiments, we set $r=64$. For the experiments, if not otherwise specified, we position the query before the documents.
@@
It is worth noting that, our method does not need to calculate another attention matrix and add the new attention matrix to the previous one. We can directly add the activation to the existing attention matrix. However, directly add the activation would greatly disrupt the attention patter, causing bad performance, therefore, we set $\xi$ to be 0 at first so our method becomes vanilla attention, then during the training process, we gradually increase $\xi$ linearly for the first 80% steps and keeps $\xi$ a constant for the last 20% steps. And besides LoRA fine-tuning, We conduct full fine-tuning under this setting and the result is shown in Table 5.
We compare our method with LoRA mainly because our method focus on how to adjust the attention schema to make the model more robust to noise. But current researchs primarily focuses on either how to structure training data (RAG_robust; RAG_robust_finetune_25) or how to train models to better handle different types of noise (RAAT). In contrast, our work addresses a different challenge: the standard self-attention mechanism inherently struggles to filter out noisy information So we modify the attention computation by incorporating a non-linear activation function to enhance robustness. Differential Transformer (differentialtransformer) also tries to adjust attention, but requires to train the model from scratch instead of fine-tuning based on existing model, and CrAM does not involve fine-tuning, so the comparison is unfair.
@@ 6.3 Main Results
Figure 3: The difference of attention score on answer tokens (mean(attn(answer))-mean(attn(other)), for clarity, we scale this gap by a factor of 1,000.
Table 1 shows the performance when faced with explicit distracting documents, we evaluate two different setting where the query is placed behind the documents or ahead the documents. The result shows that with $g(x)$, the model can better distinguish between relevant and distracting documents, showing better performance. Also, we can observe that after fine-tuning, placing the query ahead of documents still helps. The result shown in Table 1 might be high especially for 2wiki, this is mainly because we add gold documents to the context to show how can the model grab useful information from context. We also show the performance when all documents are retrieved in Table 4, it also shows that our method performs well.
@@
We also conduct experiments on Qwen/Qwen2.5-7B-Instruct and mistralai/Mistral-7B-Instruct-v0.2, we show the performance in Table 2, which shows that our method also effectively helps the performance on Qwen2.5-7B-Instruct and Mistral-7B-Instruct. We also show the performance with full fine-tuning in Table 5.
@@ 6.4 The Difference of Attention Score
By adding the activation function, our method can effectively distinguish between useful information and noise, to show this, we calculate the attention score gap between the tokens containing answer and other tokens (after softmax, mean(attn(answer))-mean(attn(other)), for clarity, we scale this gap by a factor of 1,000, the result is show in Figure 3. We do not show the result of ASQA because it is a long form QA, the answer is not directly indicated in the documents. The result shows that the rectification helps the model to recognize the answer and filter out noisy information, which explains why our method helps.
6.5 Ablation Study with Different Activation
Figure 4: The performance when we set different activation function, $sum(\xi\cdot tanh(x),x)$ (sum), $\xi\cdot tanh(x)$ (tanh). We also show the performance when we use $g(x)=max/min(x,2\xi\cdot(sigmoid(x)-0.5))$ (sigmoid). $\xi$ is set to 3, and ours means our method.
We also conduct experiments when $g(x)=\xi\cdot tanh(x)$ and $g(x)=\xi\cdot tanh(x)+x$. $g(x)=tanh(x)$ stands for the situation that $g(x)$ only focus on enlarge the margin between relevant and irrelevant tokens without further linear growth. And $g(x)=tanh(x)+x$ stands for the situation where the the steady growth process is missing, it rapidly increase with $x$, so the soft clamp will not work. We also try using
$g(x)=\begin{cases}\max(\xi\cdot(sigmoid(x)-0.5)),x)&\text{ if }\ x>=0,\\ \min(\xi\cdot(sigmoid(x)-0.5)),x)&\text{ else, }\end{cases}$
We use sigmoid activation instead of tanh and show the performance. As shown in Figure 4, simply use $tanh(x)$ or $tanh(x)+x$ has suboptimal performance, this is mainly because they fail to optimize the attention pattern on relevant tokens or missing the saturating behavior. Also we can observe that replacing tanhtanh with sigmoidsigmoid helps the performance, this is mainly because sigmoidsigmoid is actually quite similar with tanhtanh, they all increases fast at the beginning and result in a steady region after the growth.
@@ 7 Conclusion
In this paper, we highlight that noise filtering in RAG is inherently difficult, limited number of transformer layers can not effectively solve it, so we require the LLM to be robust to noise information. Then we show that simply fine-tuning the LLM may not be optimal as it will disturb the attention pattern. Then we propose a new fine-tuning method which can help to be more robust to noise, extensive experiments show that our method works well, it can effectively filter out noise while taking advantage of relevant information.
@@
The paper discusses the limitation of LLM when dealing with noisy information, showing that current LLMs can not effectively process noisy information. However, although a new fine-tuning method is proposed, it can not fully address the problem as it is a fine-tuning based on the trained model. It might help more if we train a model from scratch, but due to limited computational resource, this can only leave for future work.
@@ A.2 Placing the Query Ahead Helps
Table 4: The performance when all documents are retrieved. We retrieve top 5 documents and use those retrieved documents as context. Our method also shows better performance
NQ TriviaQA HotpotQA 2wiki ASQA mean
Llama SFT 67.8 74.6 88.7 96.4 45.7 74.6
ours 69.2 76.4 90.2 96.7 49.3 76.4
Qwen SFT 65.2 72.3 85.1 96.1 49.2 73.6
ours 67.5 75.1 87.3 96.1 52.1 75.6
Mistral SFT 65.2 73.6 86.1 96.4 50.3 74.3
ours 67.6 75.2 88.3 97.8 53.6 76.5
Table 5: The performance of full fine-tuning, and ours means adding the rectification directly to the existing attention matrix, which means we only need to calculate one attention score.
@@ Appendix B Proofs
B.1 The Triple-Wise problem
Suppose Alice and Bob are given inputs $a,b\in\{0,1\}^{n}$, respectively, with the goal of jointly computing $\mathrm{DISJ}(a,b)=\max_{i}a_{i}b_{i}$ by alternately sending a single bit message to the other party over a sequence of communication rounds. Any deterministic protocol for computing $\mathrm{DISJ}(a,b)$ requires at least $n$ rounds of communication.
$r_{i}=\begin{cases}0&\text{ if }\ \exists\ a,b\ s.t.\ g(x_{i},x_{a},x_{b})=0\\ 1&\text{ else }\end{cases}$
In normal cases, judging the value of $r_{i}$ requires calculating $g(x_{i},x_{a},x_{b})$ for all $a\in[0,n_{d})$ and $b\in[n_{d},n_{d}+n_{q})$. Here we simplify the question, and we consider the situation where $g(x_{i},x_{a},x_{b})=0$ only if $b=a+n_{d}$ and $n_{q}=n_{d}$. Apparently, this is a special case of the original problem, and if one layer of self-attention fail to solve this, it is impossible for it to solve the original problem.
If we assume that the input is like,
${\bm{x}}_{i}\in\begin{cases}\{{\bm{x}}_{i}\}&\text{if}\ i=0,\\ \{0,{\bm{x}}_{a}\}&\text{if}\ i\in\{1,\dots,n_{d}-1\},\\ \{0,{\bm{x}}_{b}\}&\text{if}\ i\in\{n_{d},\dots,2\cdot n_{d}-1\}.\end{cases}$ (2)
Given input $(a,b)\in\{0,1\}^{n_{d}}\times\{0,1\}^{n_{d}}$, let $x_{i}=x_{a}$ if and only if $a_{i}=1$ and let $x_{i}=x_{b}$ if and only if $b_{i-n_{d}}=1$. In this way $r_{i}=0$ if and only if $\mathrm{DISJ}(a,b)=1$.
For simplicity, we use $n=n_{d}$. If we consider the setting of RAG, then actually Alice and Bob each hold a matrix ${\bm{A}}\in\mathbb{R}^{n\times d}$, ${\bm{B}}\in\mathbb{R}^{n\times d}$, each row of the matrix contains $\bm{w}$}, which is the embedding to judge if the token a relevant information. So $d=H({\bm{w}})$. and we assume that ${\bm{x}}=[{\bm{s}},{\bm{w}}]$.
Then, to judge is a token relevant, we need to embedding $x_{i}$ to contain information about $\bm{w}$}, if we want to judge the relevance of $x_{0}$, then
${\bm{x}}_{i}=\begin{cases}{\bm{x}}_{i}&\text{if}\ i=0,\\ [{\bm{s}},{\bm{a}}_{i}]&\text{if}\ i\in\{1,\dots,n_{d}-1\},\\ [{\bm{s}},{\bm{b}}_{i}]&\text{if}\ i\in\{n_{d},\dots,2\cdot n_{d}-1\}.\end{cases}$ (3)
Let $\mathrm{DISJ1}({\bm{A}},{\bm{B}})=\max_{i}(g^{\prime}({\bm{x}}_{i},{\bm{a}}_{i},{\bm{b}}_{i}))$ to be noise filtering task in RAG, then it requires to access $\bm{w}$} of all tokens, which means the calculation of $DISJ1$ requires $n\times H({\bm{w}})$ bits of communication.
Also similar to the setting of $x$, $r_{i}=1$ if and only if $\mathrm{DISJ1}({\bm{A}},{\bm{B}})=1$.
Then, this is the same with the 3Match problem, following the proof of Theorem 7 in 3Match, and with the following form of transformer $f({\bm{X}})$, $2pH\log\log n+mpH\log\log n\approx mpH\log\log n$ bits are communicated.
$f({\bm{X}})=\phi(f_{h}({\bm{X}})),$
where $\phi$ stands for the feed forward layers.
$f_{h}({\bm{X}})=\frac{\sum_{i=1}^{N}\exp\left(({\bm{W}}_{q}x_{1})^{T}{\bm{W}}_{k}x_{i}\right){\bm{W}}_{v}x_{i}}{\sum_{i=1}^{N}\exp\left(({\bm{W}}_{q}x_{1})^{T}{\bm{W}}_{k}x_{i}\right)}$
Therefore, only we require $mpH\log\log n\geq nH({\bm{w}})\rightarrow mph\geq nH({\bm{w}})/\log\log n$.
Theorem B.1.
For input documents of length $n$, if $mpH\leq\Omega(nH({\bm{w}})/\log\log n)$, then there is no one layer transformer $\mathcal{M}$} with embedding size $m$, precision $p$ and $H$ heads satisfying ${\mathcal{M}}(X)=r$.
Also, as shown in 3Match, multiple layers of multi-headed attention are subject to the same impossibility
Conjecture B.2.
Every multi-layer transformer that computes Match3 must have width, depth, embedding dimension, or bit complexity at least $N^{\Omega(1)}$.
This is based on the situation that, each translation only need 1 bit, for noise filtering, we require to translate $H({\bm{w}})$ bit of information each time, so it requires width, depth, embedding dimension, or bit complexity at least $N^{\Omega(1)}\cdot H({\bm{w}})$. This directly means that triple-wise problems can not be solved with limited number of transformer layers
B.2 Proof of Theorem 4.1
@@ B.1 The Triple-Wise problem
B.3 MLP also fails to filter out noise